For the quadratic function Q(x)=x^2+5x-36, determine the exact value of the minimum.
Answer by scott8148(6628)

the minimum value (or the maximum value if x^2 was negative) lies on the axis of symmetry

the equation for the axis of symmetry is x=-b/2a ... in this case, x=-5/2

to find the value for the minimum, just plug into the equation ... Q(x)=(-5/2)^2+5(-5/2)-36 ... Q(x)=-42.25